OpenStudy (anonymous):

How can I implicitly differentiate the following... x^2 + y^2 = (2x^2 + 2y^2 - x)^2

OpenStudy (unklerhaukus):

with respect to x?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

I guess it would be chain rule on both sides but it gets a bit messy when I do it

OpenStudy (unklerhaukus):

\[x^2 + y^2 = 2x^2 + 2y^2 - x\\ \frac{d}{dx}\left(x^2 + y^2\right) = \frac{d}{dx}\left(2x^2 + 2y^2 - x\right)\\ \frac{d}{dx}(x^2)+\frac{d}{dx}(y^2)=\frac{d}{dx}(2x^2)+\frac{d}{dx}(2y^2)-\frac{d}{dx}(x)\]

OpenStudy (anonymous):

I just corrected the question, the right side is to the power of 2

OpenStudy (unklerhaukus):

Implicitly differentiate the terms with \(y\), like this: \[\frac{d}{dx}(y^2) = 2y\frac{dy}{dx}\]

OpenStudy (unklerhaukus):

oh ok, what do you get for the left hand side?

OpenStudy (anonymous):

I got 2x + 2yy'

OpenStudy (unklerhaukus):

good

OpenStudy (unklerhaukus):

now the righthand side \[=\frac d{dx}(2x^2+2y^2-x)^2\]

OpenStudy (anonymous):

Would it be 2(2x^2 + 2y^2 -x)(4x +4yy' -1) ?

OpenStudy (unklerhaukus):

looks good!

OpenStudy (anonymous):

Now the simplification XD

OpenStudy (unklerhaukus):

hmmm, i don't think it simplifies very far, implicit derivatives often don't simplify nice

OpenStudy (unklerhaukus):

What do you get?

OpenStudy (amilapsn):

please look @ my question after this @UnkleRhaukus

OpenStudy (anonymous):

This is what I get at the end ... \[\frac{ 2(4x^3 - 3x^2 +4xy^2 - y^2) }{ y(1-8x^2 - 8y^2 +4x )}\]

OpenStudy (anonymous):

@UnkleRhaukus is this what you get as well?

OpenStudy (unklerhaukus):

what happened to the equals sign?

OpenStudy (anonymous):

What I'm trying to solve is actually this question...

OpenStudy (anonymous):

Oh lol forgot that... y' = to all that ^

zepdrix (zepdrix):

Ooo I think the same derivative :) I think you did that correctly

zepdrix (zepdrix):

Gave yourself a little extra work maybe though.. hmm

zepdrix (zepdrix):

I got the same derivative* woops typo

OpenStudy (anonymous):

Oh that's a relief! Haha at least i got that right XD

OpenStudy (unklerhaukus):

... I got 2x + 2yy' = 2(2x^2 + 2y^2 -x)(4x +4yy' -1) x + yy' = (2x^2 + 2y^2 -x)(4x +4yy' -1) = +/- √(x^2 + y^2)(4x +4yy' -1) how do you simplify it?

OpenStudy (anonymous):

Well I only solved for y'

zepdrix (zepdrix):

It probably would have been easier to simply plug in your coordinate from this point,\[\Large\rm 2x+2yy'=2(2x^2+2y^2-x)(4x+4yy'-1)\]And simply add all the stuff up before solving for y'. Especially since you're getting a bunch of 0's in there.

OpenStudy (anonymous):

Ohhh right!!! XD But then Im gonna be left with y in the equation as well, how would I plug in the slope in the point-slope formula at the end for finding the equation of the tangent line?

zepdrix (zepdrix):

Left with y in the equation? No no, you're plugging 0s in for the Xs, and 1/2s in for the Ys, ya? :) Only #s and `Y'`s left over.

zepdrix (zepdrix):

y' = m <-- this i the slope you're looking for. We would STILL have to do the work to solve for y' to find our slope, but im simply saying that if you plug in the coordinate BEFORE solving for y', it's a ton easier, because you're plugging in a bunch of 0s.

OpenStudy (anonymous):

Okay - so we plug in x=0 and y=1/2 and then solve for y'? Then plug that number in for 'm' in the point-slope formula, correct? :D

zepdrix (zepdrix):

\[\Large\rm y-\frac{1}{2}=m(x-0)\]Yes good :) Figure out that m! And since our line is tangent to the curve, it "touches" right at this coordinate point, so both the curve and this tangent line share the point, we can use it for our point-slope form.

OpenStudy (anonymous):

Alright so I'm gonna give that a try and tell you what I get :)

OpenStudy (anonymous):

Please don't leave! :D

zepdrix (zepdrix):

fine fine fine -_-

OpenStudy (anonymous):

Lol so plugging in the x and y values I got -1... Tell me I'm not wrong :D

zepdrix (zepdrix):

Let's look at the graph to sort of check our work easily.

zepdrix (zepdrix):

Here is the point (0,1/2)